Which statement is true of x^1/3 and -x^1/3?
The relationship cannot be determined from the given information.
The quantity in -x^1/3 is greater.
The two quantities are equal.
: Which statement is true of x^1/3 and -x^1/3?
The relationship cannot be determined from the given information.
The quantity in -x^1/3 is greater.
The two quantities are equal.

Answer by midwood_trail(310)

Which statement is true between x^1/3 and -(x^1/3)?
The relationship cannot be determined from the given information.
The quantity in -x^1/3 is greater.
The two quantities are equal.
Hum...x^(1/3) = cube root of {x} and -(x^1/3) = negative cube root of {x}.
If we let x = 64 in both cases as an example, we have this:
64^(1/3) = cube root of {64} = 4.
-(64^1/3) = negative cube root of -{64} = -4.
As you know, 4 is > than -4. So, my conclusion is that both quantities are not equal and that x^(1/3) is greater than -(x^1/3).
Is this what you are looking for?